Determine the complex number z which satisfies the equations |z + 3i| = |z + 5 − 2i| and |z − 4i| = |z + 2i| simultaneously
1
Expert's answer
2020-05-20T17:03:56-0400

"z=x+iy"

Hence

"|z+3i|=\sqrt{x^2+((y+3)^2)};|z+5-2i|=\sqrt{(x+5)^2+((y-2)^2)}"

"\sqrt{x^2+((y+3)^2)}=\sqrt{(x+5)^2+((y-2)^2)}"

"x^2+(y+3)^2=(x+5)^2+(y-2)^2"

From the other equality we obtain

"|z-4i|=\sqrt{x^2+(y-4)^2};|z+2i|=\sqrt{x^2+(y+2)^2}"

"x^2+(y-4)^2=x^2+(y+2)^2"

Therefore

"(y-4)^2=(y+2)^2"

"-8y+16=4y+4"

"12y=12"

"y=1"

"x^2+16=(x+5)^2+1"

"10x+26=16"

"x=-1"

"z=-1+i"